FTIR SPECTROSCOPY OF THE ν_{4} BANDS OF ^{14}NO_{3} and ^{15}NO_{3}

R. FUJIMORI, N. SHIMIZU, J. TANG, \underline{K. KAWAGUCHI}, Department of Chemistry, Faculty of Science, Okayama University, 3-1-1, Tsushima-Naka, Okayama 700-8530, Japan; T. ISHIWATA, Graduate School of Information Sciences, Hiroshima City University, 3-4-1 Otsuka-Higashi, Hiroshima 731-3194, Japan,.

High-resolution Fourier transform infrared spectra of the NO_{3} ν_{4} bands of ^{14}N and ^{15}N isotopic species were observed in the 365 and 360 cm^{−1} region. Observed lines were used to determine ΔK=3 combination differences of the ground state coupled with transition frequencies of the ν_{3} hot band ν_{3} + ν_{4} - ν_{4} in 1127 cm^{−1} region^{a} and ν_{3} + ν_{4} band in 1492 cm^{−1} region, which made it possible to determine the rotational C_{0} constants of 0.228 6321(67) cm^{−1} and 0.228 674(11) cm^{−1} for ^{14}NO_{3} and ^{15}NO_{3} , respectively. Inertial defects of several vibrational states were calculated by using a formula of Jagod and Oka^{b} and compared with the observed values. Although the observed Coriolis coupling constants ζ_{4}=-0.188 (^{14}NO_{3} ) and -0.156 (^{15}NO_{3} ) are largely different from the expected value of -0.7 from a force field calculation, observed inertial detects are in good agreement with calculated values, for example, Δ_{obs}(Gr.)= 0.206 [0.223] amuÅ^{2}, (v_{4}=1) 0.434[0.437] amuÅ^{2} etc, where calculated values are given in square brackets.

aK. Kawaguchi, N. Shimizu, R. Fujimori, J. Tang, T. Ishiwata, and I. Tanaka, J. Mol. Spectrosc. 268, 85 (2011).

bM.-F. Jagod and T. Oka, J. Mol. Spectrosc. 139, 313 (1990).